You are analyzing the following two mutually exclusive projects and have developed the following information. Please calculate the IRRs for the two projects and the crossover rate. (6 points) Which project should you accept if the cost of capital is 5%, and which project should you accept if the cost of capital is 10%? (4 points) 


year project A project B

0 -1100 -900

1 400 700

2 500 100

3 100 100

4 400 200

IRR A: ___________ IRR B: ___________ Crossover Rate: ____________ If WACC=5%, accept _______ If WACC=10%, accept _______
Answer #1

Using financial calculator to calculate
CF0=-1100;CF1=400;CF2=500;CF3=100;CF4=400;CPT IRR =11.13%

Using financial calculator to calculate
CF0=-900;CF1=700;CF2=100;CF3=100;CF4=200;CPT IRR =12.22%

Incremental Cash Flow
CF0=-1100+900=-200
CF1 =400-700 =-300
CF2=500-100=400
CF3=100-100=0
CF4=400-200=200
CF0=-200;CF1=-300;CF2=400;CF3=0;CF4=200;CPT IRR =9.36%

If WACC =5%
NPV of Project A =-1100+400/1.05+500/1.05^2+100/1.05^3+400/1.05^4=149.93
NPV of Project B =-900+700/1.05+100/1.05^2+100/1.05^3+200/1.05^4=108.29
At 5% Accept Project A

If WACC =10%
NPV of Project A =-1100+400/1.10+500/1.10^2+100/1.10^3+400/1.10^4=25.20
NPV of Project B =-900+700/1.10+100/1.10^2+100/1.10^3+200/1.10^4=30.74
At 10% Accept Project B